OpenStudy (anonymous):

Miguel is 46years old. he is 4 years older than thrice his son's age. Find the age of his son.

OpenStudy (mathteacher1729):

Were you able to translate this word problem into an equation?

OpenStudy (saifoo.khan):

(3x+4) = 46 If im not wrong.

OpenStudy (mathteacher1729):

Because there are two parts to this prob: 1_ Translate words to equation 2_ Solve equation and find the desired age

jhonyy9 (jhonyy9):

so x=42/3 =14